OpenStudy (anonymous):

If the sequence is geometric, what is the common ratio? 500, 100, 20, 4,... A. 5 B. 125 C. 1/5 D. not geometric

OpenStudy (anonymous):

500 / 100 = 100 / 20 = 20 / 4 = 5